**not really sorta a number line question** what is the mean of the integers from -410 to 400. Please explain
Found 2 solutions by CubeyThePenguin, MathTherapy:
Answer by CubeyThePenguin(3113)

sum from -410 to 400: -1 - 2 ... -10 = -55

number of numbers: 400 - (-410) + 1 = 811

mean = -55/811

**My answer is correct. This questions for the mean of the INTEGERS FROM -410 to 400, not the integers -410 and 400 themselves.

-410 - 409 - 408 ... + 399 + 400 = (-400 - 399 ... - 1) + (1 + 2 + ... + 400) + (-10 - 9 - ... - 1)

The sum of the first 400 numbers cancel out, leaving only the sum from -1 to -10, which is -55.

You are right for the second part, however. I made a small calculation mistake. There are 811 numbers from -410 to 400, not 821. Next time, I would really appreciate your solution so I can understand my mistakes as well.

Answer by MathTherapy(10557)

**not really sorta a number line question** what is the mean of the integers from -410 to 400. Please explain
Don't know where on EARTH the other person gets some of his answers. Mean is SIMPLY: 
The SUM of the INTEGERS from - 410 to + 400 is CERTAINLY NOT - 55, and the number of INTEGERS in that range is CERTAINLY NOT 821 either!!
I need NOT say ANYMORE!!!